### v2

# Chirality-Driven Magnetization Emerges from Relativistic Four-Current Dynamics

Abstract
Chirality-induced spin selectivity (CISS) is a striking quantum phenomenon in which electron transport through chiral molecules leads to spin polarization—even in the absence of external magnetic fields or magnetic components. Although observed in systems such as DNA, helicenes, proteins, and polymers, the fundamental physical origin of CISS remains unresolved. Here, we introduce a time-dependent relativistic four-current framework, in which charge and current densities evolve according to the time-dependent variational principle. Real-time relativistic four-current simulations enable direct analysis of helical currents and induced magnetization dynamics. Applied to helicenes—axially chiral molecules lacking stereocenters—our simulations reveal curvature-induced helical electron currents that generate spontaneous magnetic fields aligned along the molecular axis. These fields are handedness-dependent and reach magnitudes of $10^{-1}$ Tesla per single helicene strand. Our results suggest that CISS may arise from intrinsic, relativistic curvature-induced helical currents and the associated magnetic fields within chiral molecules. This four-current mechanism offers a self-contained explanation for the driving force underlying spin selectivity, independent of interfacial effects or unphysically enhanced spin–orbit coupling. Furthermore, our results provide a new perspective that offers a unifying framework with the potential to reconcile many existing hypotheses and theoretical models, while also suggesting several testable predictions that can be examined experimentally.
[…]
Chiral molecules play a vital role in biological processes,1, 2, 3, 4, 5, 6, 7, 8 and have emerged as promising platforms for engineering spin-dependent technologies.9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21 Owing to their lack of inversion symmetry, these molecules exhibit a strong coupling to electron spin, manifesting in the phenomenon known as chirality-induced spin selectivity (CISS).22, 23 The CISS effect was first observed in DNA,24, 25, 26 where electron transport through chiral structures led to spin-polarized currents even in the absence of an external magnetic field.
Since this pioneering discovery, the CISS effect has been experimentally confirmed in a broad range of systems, including helicenes,27, 28, 29, 30, 31, 32, 33, 34 proteins,35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46 and synthetic polymers.47, 48, 42, 49, 50, 51, 52 Although numerous theoretical models have been proposed to explain the underlying mechanism,53, 54, 55, 56, 57, 58 the physical origin of the CISS effect remains a subject of ongoing debate and investigation.
Contemporary theoretical explanations for the CISS effect can be broadly categorized into two classes. The first viewpoint posits that an effective spin–orbit coupling term must be added to the Hamiltonian describing the electronic structure of a chiral system. However, such a framework struggles to account for the observation of CISS effects in certain hydrocarbon systems, where intrinsic spin-orbit coupling is negligibly small.59, 60
The second viewpoint attributes spin selectivity primarily to interface effects between electrodes or substrates and the chiral molecules. In this class of theories, spin–orbit interactions in the substrate convert the orbital angular momentum, filtered by the chiral molecule into spin angular momentum, resulting in spin-polarized transmission.61, 58, 62, 63 While this mechanism offers a plausible explanation for surface-bound systems, it fails to explain recent gas-phase experiments,64 where clear CISS effects were observed in the complete absence of electrodes or supporting substrates.
These inconsistencies highlight the need for a more general and intrinsic explanation of the CISS phenomenon—one that captures the fundamental role of chirality in shaping spin dynamics, independent of external interfaces or large spin–orbit couplings. In this context, a fully ab initio relativistic electron dynamics based on the four-current formalism may offer a promising path forward.
Accurately describing electron dynamics in atoms, molecules, and materials remains a central challenge in modern theoretical chemistry and physics. Real-time time-dependent density functional theory (RT-TDDFT)65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77 and time-dependent wave function approaches, such as time-dependent coupled-cluster (TD-CC)78, 79, 80, 81, 82, 83, 84, 85, 86 and time-dependent configuration interaction (TD-CI)87, 88, 89, 90, 91, 92, have enabled practical simulations of electronic excitations and real-time electron dynamics.93, 94 However, conventional formulations of these methods often neglect essential relativistic and gauge-covariant aspects critical to the accurate modeling of current-carrying systems. A fully consistent and complete description of electron dynamics requires adopting the four-current formalism, which includes both the charge density and the three-component current density vector.
The current density is the most important quantum mechanical property that underlies the magnetic response and dynamics of the electrons, including NMR spectroscopy, magnetic susceptibilities, spin transport, and magnetically induced current pathways in molecules and materials.95, 96, 97, 98, 99 Within the Schrödinger framework, current can be introduced as a velocity density.100, 101 However, the nonrelativistic velocity density fails to capture spin-dependent currents or those arising from spin-couplings (e.g., spin–orbit, spin–spin), such as those described by the Dirac–Coulomb–Breit operator.102, 103, 104 The full treatment of the four-current within the four-component relativistic framework was developed by Saue and co-workers for studies of magnetically induced current density,105 and was later extended to the London orbital framework.106
Figure 1: Simulation workflow. The relativistic two-component TDDFT equations are propagated in real time, generating the four-current density on-the-fly via the Gordon decomposition. Integration of the resulting four-current reveals a spontaneous magnetization arising from chiral, curvature-induced helical currents.
In this work, we present a time-dependent relativistic four-current framework in which both charge and current densities evolve according to the time-dependent variational principle. Figure 1 illustrates the workflow of the relativistic four-current dynamics approach. The simulation begins with relativistic two-component RT-TDDFT electron dynamics. The resulting two-component density is then transformed into the four-current formalism using the Gordon decomposition. Spatial integration of the four-current yields an effective magnetic field.
Figure 2: Computed ECD spectra of (L)-[24]helicene and (R)-[24]helicene.
Figure 3: Highest singly occupied molecular orbital of anionic [24]helicene: (top) field-free ground state and (bottom) under a 0.01 au electric field aligned along the z-z-direction.
We apply this framework to investigate CISS effects in a helicene system using real-time relativistic exact-two-component RT-X2C-TDDFT quantum dynamics. Although helicenes do not possess chiral carbon stereocenters, they exhibit axial chirality, which plays a crucial role in determining their electronic dynamics. As illustrated in Fig. 2, the computed electronic circular dichroism (ECD) spectra confirm the chiral nature of L- and R-[24]helicene.
Figure 4: Time-resolved charge density, shown as the difference $\delta\rho_{\text{s}}=\rho_{\text{s}}(t)-\rho_{\text{s}}(0)$, at selected time points for (L)-[24]helicene and (R)-[24]helicene.
Specifically, we aim to simulate the experimental condition in which a helical system is subjected to an electric field bias, resulting in electron transport through the helicene.27, 29, 30, 31, 32, 33, 34 For the electron transport simulations, an extra electron is added to the helicene molecule, forming an anionic molecular system. The introduction of a spin-1/2 system with an extra electron is specifically designed to probe the spin-current contribution, which can be rigorously described within the relativistic four-current formalism.
The spatial distribution of this additional electron in the ground state is shown in Fig. 3. In the absence of an external electric field, the electron is delocalized across the entire molecule, with the highest density concentrated near its center. The relativistic RT-X2C-TDDFT electron dynamics are initiated from an electric-field-polarized state (see Theory and Computational Details for more information), with the initial molecular orbital depicted in Fig. 3. The quantum dynamics are driven by an external field $E_{z}=-0.01$ au in the z-z direction. The total simulation time is 1 fs, and a complex absorbing boundary condition was applied.107 During the simulation, the electrons did not reach the boundaries; therefore, no reflection occurred, and we observed no significant difference between the scenarios with and without the complex absorbing boundary condition. This setup represents a non-equilibrium transport condition.108, 109, 107
Relativistic four-current dynamics simultaneously capture the time-dependent behavior of both charge $\rho_{\text{s}}$ and current ($J_{x}$, $J_{y}$, $J_{z}$) densities. We begin by examining the evolution of the charge density. Figure 4 presents the time-resolved charge or scalar density, shown as the difference density $\delta\rho_{\text{s}}=\rho_{\text{s}}(t)-\rho_{\text{s}}(0)$, at selected time points. While the charge density evolution reflects a typical electron transfer process, the dynamics indicate that the electron density remains confined to the helicene backbone. This suggests that the helical curvature of the structure plays a key role in shaping the characteristics of charge transport.
The helical curvature of the molecule can induce a rotational component in the electron current. This effect is demonstrated in Fig. 5, where the vector form of the electron current, represented by (${J}_{x}(\mathbf{r},t)$, ${J}_{y}(\mathbf{r},t)$, ${J}_{z}(\mathbf{r},t)$), is shown as a function of space at $t=0.1$ fs when the helical-current is at its maximum at the center of the helicene. Rather than exhibiting a simple axial velocity current along the $z$-direction, the current vector displays a pronounced rotational behavior that tracks the helical curvature of the structure, as seen in Fig. 5. This rotation highlights the direct influence of molecular chirality on current dynamics.
Figure 5: Side and top views of the helical current in (L)-[24]helicene and (R)-[24]helicene at 0.10 fs.
To examine the current vector in more detail, Fig. 5 displays the current direction in the plane perpendicular to the chiral axis (the $z$-direction). As shown, (L)-[24]helicene produces a counterclockwise current, whereas (R)-[24]helicene generates a clockwise current when viewed from above, looking down along the z-z-axis. We would like to emphasize that in both simulations, the direction of charge transfer is identical, specifically, from +z+z to z-z. This suggests that the observed difference in current circulation arises solely from the molecular handedness.
Figure 6: Side and top views of the helical-current-induced magnetic field in (L)-[24]helicene and (R)-[24]helicene at 0.10 fs.
According to classical electrodynamics, helical currents within a molecule generate a magnetic field whose direction can be determined using the right-hand rule. In the case of (L)-[24]helicene and (R)-[24]helicene, the handedness of the molecule causes the induced magnetic fields to point in opposite directions. Within the ab initio framework, this emergent magnetic field can be computed by integrating the current density using Jefimenko’s equations (see the Theory and Computational Details section for details). Figure 6 illustrates the emergence of a magnetic field induced by the progression of helical curvature-induced currents. Notably, (L)-[24]helicene and (R)-[24]helicene produce magnetic fields of similar magnitude but opposite direction, reflecting their mirror-image structures.
To investigate the length dependence of the helical-current-induced magnetic field, a series of helicenes with varying heights were simulated using four-current dynamics. The maximum magnetic field generated for each helicene is shown in Fig. 7. The field strength is $\sim 10^{-1}$ Tesla. As shown in the figure, the magnetic field increases with helicene height and reaches a plateau beyond [16]helicene. This trend suggests the presence of a critical helical length beyond which the helical-current-induced magnetic field saturates. This behavior can be attributed to the finite delocalization length of $\pi$-conjugated electrons, which governs low-energy electron transport in such systems.
Figure 7: The maximum generated helical-current-induced magnetic field for a variety of helicene lengths.
The emergence of the helical-current-induced magnetic field is a likely physical driving force underlying the CISS effect. A magnetic field can break spin (Kramers’) symmetry, lifting the degeneracy between spin-up and spin-down states (see Fig. 8 for an illustration). The direction of the field determines which spin manifold is energetically favored, thereby enabling a spin-filtering effect.
From the Boltzmann population analysis of the Kramers’ pair (spin-down and spin-up) at 5 K under a 0.46 T magnetic field generated by the helical current in [24]helicene, the spin-down/spin-up ratio is approximately 0.9, indicating that the magnetic field favors one spin manifold by about 10%. However, this small magnitude suggests that the helical-current-induced magnetic field, although able to break time-reversal symmetry, is by itself insufficient to account for the full extent of the CISS effect. Additional mechanisms must therefore exist to amplify the magnetic field induced by the helical current.
Figure 8: A schematic representation of the helical-current-induced magnetic field giving rise to broken spin (Kramers’) symmetry.
Figure 9: Contributions of paramagnetic, diamagnetic, and spin-dependent paramagnetic terms to the magnetic field induced by helical currents. The inset highlights the spin-dependent paramagnetic contribution. The helicene structure is shown with current arrows indicating the point where the time-dependent magnetic field is evaluated.
The four-current expression comprises three terms, each offering key insights into the nature of the helical current. As shown in Fig. 9, the velocity current (first term in Eq. 5) dominates, being approximately two to three orders of magnitude larger than the spin-dependent paramagnetic current (third term). The diamagnetic term (second term) is absent due to the lack of an external vector potential. Both the velocity and spin-dependent currents originate from the molecular helical curvature; the latter additionally depends on the intrinsic angular momentum of the system.
From the perspective of a foundational theoretical framework, several important conclusions can be drawn. The velocity current (first term) can be fully accounted for within non-relativistic quantum dynamics, indicating that the observed spontaneous generation of the magnetic field has a non-relativistic origin. Although the spin-dependent paramagnetic current (third term) requires a fully variational relativistic treatment, its contribution in the helicene system is negligible, indicating that a self-consistent treatment of spin and spin–orbit interactions is not necessary in this case. However, in systems with strong spin and spin–orbit couplings, the spin-dependent paramagnetic current can become significant, demanding the use of relativistic four-current dynamics. A fully self-consistent treatment of the diamagnetic term (second term) requires a quantum field description, which could be approximated using a semiclassical approach in future work.
Conclusion
The physical origin of chirality-induced spin selectivity (CISS) effects in helicenes was investigated using relativistic four-current dynamics within the RT-X2C-TDDFT framework. Simulations revealed a curvature-induced helical current arising from the electron dynamics in the helicene structure. Integration of this helical current over the transport region generates a spontaneous magnetic field aligned with the helical axis. Although the direction of electron transport remains the same in both L- and R-helicenes, the direction of the resulting helical-current-induced magnetic field depends on the molecular handedness, providing compelling evidence for a chirality-induced magnetic response.
The computed helical-current-induced magnetic field for a single helicene lies in the range of $10^{-1}$ Tesla. Its magnitude increases with the length of the helicene, reaching a maximum at a critical threshold length, around [16]helicene, beyond which further increases in length do not significantly enhance the field.
This work examines the origin of spontaneous magnetization arising from helicity-induced electron currents. As shown in Eq. 6, if the four-current does not have a preferential axis or is not associated with a principal axis of electron transport, the isotropic averaging of the induced field vanishes, resulting in no net magnetization. In contrast, when electron transport occurs along a well-defined direction in a helical structure, the resulting preferential alignment of the four-current breaks isotropic cancellation and gives rise to a non-vanishing spontaneous magnetization. This highlights the fundamental connection between helicity, directional electron motion, and the breaking of isotropic symmetry, which together underpin the emergence of magnetization in chiral systems. The helical-current-induced magnetic field is likely the origin of the CISS effect by breaking spin symmetry and favoring one spin orientation, enabling spin filtering.
While the system studied here is specifically tied to helicity-induced electron currents, the four-current quantum dynamics framework is completely general. It enables the analysis of a broad range of driving forces in diverse physical and chemical contexts, including systems with intrinsic chiral centers, strong spin–orbit interactions, and other symmetry-breaking mechanisms.
Perspective and Hypotheses
The physical insights gained from this work provide a new perspective on the CISS effect. While the helical-current-induced magnetization accounts for spin Zeeman splitting, it must be amplified to explain the full extent of spin selectivity. This consideration introduces the concepts of a ‘driving force’ and ‘amplification’ in CISS, offering a unifying framework with the potential to reconcile many of the existing hypotheses and theoretical models. The ‘driving force’ is the helical current, which may arise from electron transport through a helical structure and/or from chiral phonons, the latter also correlating with temperature effects. Potential ‘amplification’ mechanisms include strong spin–orbit coupling, which enhances spin-dependent paramagnetic contributions; molecular aggregation, which increases the overall current; and quantum-field self-consistent interactions, which strengthen the diamagnetic response.
Together, these analyses and simulation results form the foundation for informed hypotheses regarding the physical origin of the CISS effect:
  • The low-energy conduction band electrons in organic helical systems possess a finite delocalization length. Once this threshold is exceeded, additional $\pi$-conjugation no longer contributes effectively to the helical-current-induced magnetic field.
  • The helical pitch and radius determine the curvature-induced helical current. Therefore, molecular engineering of these structural parameters offers a viable route to modulate the magnitude of the helical current.
  • Although a single helical strand can generate a chirality-dependent magnetic field in the range of $10^{-2}$ to $10^{-1}$ Tesla, this field is likely too weak to fully account for complete spin filtering. Potential amplification mechanisms include strong spin–orbit coupling within the molecule or at the electrode interface, as well as the formation of ordered aggregates to enhance current density.
  • The helical-current-induced magnetic field can, in turn, enter the expression for the helical current itself as a vector potential in the second term of Eq. 5. Consequently, a self-consistent treatment of the field–current interaction may further amplify the observed magnetic field. Ideally, these effects would be rigorously treated within a quantum field framework, providing motivation for the development of four-current quantum dynamics coupled with quantum-field-enabled time-dependent electronic structure theory.
  • In insulators, mechanistic driving forces, such as chiral phonons with non-zero momentum, can transport four-current densities through a helical structure, thereby generating helical currents and spontaneous magnetic field.
Methods
Theory and Computational Details
In this paper, the following notation is adopted, unless otherwise specified:
  • $\{\mu,\nu,\lambda,\sigma\}$ : atomic orbitals (AOs)
  • $\varphi_{i}(\mathbf{r},t),i\in{p,q,r,s}$ : molecular orbitals (MOs)
Relativistic RT-X2C-TDDFT
In this work, we use the exact-two-component (X2C) transformation to yield an electron-only relativistic Hamiltonian.110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 75, 123, 124, 125, 126, 127, 128, 129 Specifically, we use the one-electron X2C transformation, which allows for one-step approach to construct a unitary transformation matrix to “fold” the small component coefficients into the large component. The one-electron X2C approach makes use of the effective one-electron spin–orbit Hamiltonian and avoids the four-component self-consistent-field procedure. In this work, we use the new Dirac–Coulomb–Breit-parameterized effective one-electron spin–orbit Hamiltonian in the X2C approach.130 The relativistic Hamiltonian does not include explicit quantum field operators; therefore, the four-current formalism employed in this work should be regarded as an effective quantum electrodynamics treatment. It is important to note that the perturbative Pauli Hamiltonian is not bounded from below and, therefore, is unsuitable for describing relativistic four-current dynamics.
The iterative solution of the two-component self-consistent-field equations yields a set of complex-valued molecular spinors,
$\psi_{p}(\mathbf{x},t)=\begin{pmatrix}\varphi_{p}^{\alpha}(\mathbf{r},t)\\ \varphi_{p}^{\beta}(\mathbf{r},t)\\ \end{pmatrix}=\begin{pmatrix}\sum_{\mu}C_{\mu p}^{\alpha}(t)\chi_{\mu}(\mathbf{r})\\ \sum_{\mu}C_{\mu p}^{\beta}(t)\chi_{\mu}(\mathbf{r})\end{pmatrix}$, (1)
where $\mathbf{x}$ is a collective coordinate of both spatial $\{\mathbf{r}\}$ and spin coordinates $\{\alpha,\beta\}$. The spatial functions $\varphi_{p}^{\{\alpha,\beta\}}(\mathbf{r},t)$ are a linear combination of real-valued atomic orbitals. $C_{\mu p}(t)$ are the time-dependent coefficients.
Time-dependent density functional theory (TDDFT) can be formulated and evolved using the Liouville von Neumann equation in matrix form,93, 94
$i\frac{\partial\mathbf{P}(t)}{\partial t}=[\mathbf{F}(t),\mathbf{P}(t)]$ (2)
where $\mathbf{F}$ and $\mathbf{P}$ are the Fock/Kohn–Sham and the density matrices in orthonormal basis.
For two-component real-time methods, the density matrix can be represented in spin-blocked form or as a sum of the scalar density $\mathbf{P}_{s}$ and the vector magnetization density $\mathbf{P}_{x},\mathbf{P}_{y},\mathbf{P}_{z}$:
$\mathbf{P}$ $=\begin{pmatrix}\mathbf{P}^{\alpha\alpha}&\mathbf{P}^{\alpha\beta}\\ \mathbf{P}^{\beta\alpha}&\mathbf{P}^{\beta\beta}\end{pmatrix}$ (3)
$=\mathbf{P}_{s}\otimes\mathbf{I}_{2}+\sum_{k\in\{x,y,z\}}\mathbf{P}_{k}\otimes\boldsymbol{\sigma}_{k}$ (4)
$\mathbf{P}_{s}$ $=\frac{1}{2}(\mathbf{P}^{\alpha\alpha}+\mathbf{P}^{\beta\beta})$
$\mathbf{P}_{x}$ $=\frac{1}{2}(\mathbf{P}^{\alpha\beta}+\mathbf{P}^{\beta\alpha})$
$\mathbf{P}_{y}$ $=\frac{i}{2}(\mathbf{P}^{\alpha\beta}-\mathbf{P}^{\beta\alpha})$
$\mathbf{P}_{z}$ $=\frac{1}{2}(\mathbf{P}^{\alpha\alpha}-\mathbf{P}^{\beta\beta})$
where $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of Pauli matrices.
Commonly used density functional approximations that only depend on the spin-collinear densities ($\mathbf{P}^{\alpha\alpha}$ and $\mathbf{P}^{\beta\beta}$) must be modified to be compatible with two-component Hamiltonians.131, 132, 133, 134, 135, 136, 114, 137, 138, 139, 115, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 124 The non-collinear DFT framework used in this work redefines the functionals (in the similar spirit of spin-polarized DFT) to depend on a set of auxiliary generalized variables which take the full magnetization vector density into account,141, 144, 145, 124 As this approach has been developed and calibrated, we refer readers to Ref. 124 for mathematical derivations and implementations and Ref. 150 for a recent review of non-collinear DFT methods.
Two-Component Four-Current
While the two-component density is convenient for implementing the time-dependent variational principle, its corresponding four-current representation offers a more physically meaningful description in capturing the interplay between charge, current, and spin dynamics, particularly in relativistic and magnetically active systems.
The Gordon decomposition of the two-component current density can be derived in the following manner. The current density is given by151
$\mathbf{J}(\mathbf{r})$ $=\frac{1}{m_{e}}\mathfrak{R}\left\{\psi^{\dagger}\boldsymbol{\sigma}(\boldsymbol{\sigma}\cdot\boldsymbol{\pi})\psi\right\}$
where the mechanical momentum $\boldsymbol{\pi}=\mathbf{p}-q\mathbf{A}$ includes the linear momentum $\mathbf{p}=-i\boldsymbol{\nabla}$ and the vector potential $\mathbf{A}=\frac{1}{2}\mathbf{B}\times\mathbf{r}$ for electrons and $\varepsilon_{abc}$ is the standard Levi-Civita tensor. Using the Dirac identity, followed by mathematical derivations and rearrangement (see SI for details), we arrive at the following working expression for computing the ab initio current:
$\mathbf{J}(\mathbf{r})$ $=\frac{1}{m_{e}}\mathfrak{I}\left\{\psi^{\dagger}\boldsymbol{\nabla}\psi\right\}-\frac{q}{m_{e}}\mathfrak{R}\left\{\psi^{\dagger}\mathbf{A}\psi\right\}$
$+\frac{1}{2m_{e}}\mathfrak{R}\left\{\left(\boldsymbol{\nabla}\times(\psi^{\dagger}\boldsymbol{\sigma}\psi)\right)\right\}$ (5)
The three terms in Eq. 5 are often referred to as the paramagnetic, the diamagnetic, and the spin-dependent paramagnetic contributions to the current density, respectively. Although the last two terms are fully real, we retain the explicit notation to provide clarity when presenting the expressions in the Pauli representation. The programmable expressions in the atomic orbital basis are provided in the SI.
Four-Current-Induced Magnetic Field
The relativistic magnetic field arising from a four-current density can be obtained from Jefimenko’s equation of the magnetic field, 152, 153
$\mathbf{B}(\mathbf{r},t_{r})$ $=-\frac{\mu_{0}}{4\pi}\int\int\int\Bigg{[}\frac{(\mathbf{r}-\mathbf{r}^{\prime})}{|\mathbf{r}-\mathbf{r}^{\prime}|^{3}}\times\mathbf{J}(\mathbf{r}^{\prime},t_{r})$
$+\frac{(\mathbf{r}-\mathbf{r}^{\prime})}{|\mathbf{r}-\mathbf{r}^{\prime}|^{2}}\times\frac{1}{c}\frac{\partial\mathbf{J}(\mathbf{r}^{\prime},t_{r})}{\partial t}\Bigg{]}d^{3}\mathbf{r}^{\prime}$ (6)
In Eq. 6, the retarded time, $t_{r}$, is relativistic phenomena arising from the finite speed of light,
$t_{r}=t-\frac{|\mathbf{r}-\mathbf{r}^{\prime}|}{c}.$ (7)
In this work, the relevant velocities are much smaller than the speed of light; therefore, it is safe to neglect relativistic retardation effects. In the physical limit of a slowly varying current density, where its time dependence can be ignored, Eq. 6 reduces to the familiar Biot–Savart law.
Computational Methodology
We constructed a periodic single helicene chain in an orthogonal lattice with lattice constants $a=20$ Å  and $b=20$ Å  to eliminate interchain interactions along the $x$- and $y$-directions. The helical axis is along the $z$-direction. The lattice constant along $z$ was optimized using the Vienna Ab initio Simulation Package (VASP).154 The electron–ion interaction was treated using the projector augmented-wave (PAW) method155 with a plane-wave energy cutoff of 525 eV. For the exchange–correlation functional, the Perdew–Burke–Ernzerhof (PBE) form of the generalized gradient approximation (GGA)156 was used, along with Grimme’s DFT-D3 dispersion correction.157 A $1\times 1\times 10$ Gamma-centered Monkhorst–Pack $k$- point mesh was employed. The optimized lattice constants were $a=20$ Å, $b=20$ Å, $c=3.515$ Å  for both left- and right-handed helicenes.
The helicene molecule was cleaved from the periodic model and further geometry-optimized using the Gaussian 16 package with the B3LYP exchange–correlation functional and Def2-SV(P) basis set. Grimme’s DFT-D3 dispersion correction with Becke–Johnson damping (DFT-D3BJ) was also included.158 The optimized helicene exhibited no imaginary frequencies, confirming a stable minimum.
To investigate the electronic dynamics, we employed the relativistic RT-X2C-TDDFT approach using a development version of the open-source Chronus Quantum package.159, 160 The initial state ($t=0$) was prepared via a self-consistent field (SCF) calculation under a static electric field of 0.01 au along the z-z-direction. To initialize the electronic dynamics, an electric field of equal magnitude but opposite direction was applied, and the system was propagated in real time using the RT-X2C-TDDFT method. The time evolution was carried out for 3 fs with a time step of 0.05 au At each step, the time-dependent two-component density was Gordon-decomposed into four-current. All real time calculations used the 6-311G(d,p) basis set.
1 Code Availability
The real-time X2C-TDDFT framework is implemented in the public version of the Chronus Quantum software package.159 All calculations were performed using executables compiled with GCC version 13.2.
2 Data Availability
[…]
3 Author Contributions
S. Upadhyay, X. Zheng, and X. Li conceived the project and developed the computer code. S. Upadhyay, X. Zheng, T. Wang, and A. Shayit carried out the calculations and analyzed the data. J. Liu, D. Sun, and X. Li acquired research funding. S. Upadhyay, X. Zheng, and X. Li wrote the manuscript with input from all authors. All authors discussed the results and approved the final version of the manuscript.
Acknowledgement
The study of the CISS effect is supported by the Air Force Office of Scientific Research, Multidisciplinary University Research Initiatives (MURI) Program under award number FA9550-23-1-0311. The study of spin-dependent chemical processes is supported by the Air Force Office of Scientific Research (Award No. FA9550-21-1-0344 to XL). The development of relativistic time-dependent method is supported by the Department of Energy in the Computational and Theoretical Chemistry program (Grant No. DE-SC0006863). The Chronus Quantum computational software development is supported by the Office of Advanced Cyberinfrastructure, National Science Foundation (Grant No. OAC-2103717).